# Multimodal Evaluation of Russian-language Architectures

### v2

@@ C.3.4 Quantitative reasoning
Static counting I A V.
Temporal counting A V.
@@
Mathematical reasoning I A V.
C.3.5 Other reasoning
Other reasoning
Other reasoning
Critical thinking I A V.
Counterfactual robustness I A V.
Problem decomposition I A V.
Comparative reasoning I A V.
Appendix D Data leakage details
@@
D.2 Neighbor Generation and Feature Extraction.
For each original data point $(t,m)$ we generate $K=24$ perturbed "neighbors". We apply four distinct perturbation techniques (masking and predicting the masks with Fred-T5 model [^20] https://hf.co/ai-forever/FRED-T5-1.7B, zmitrovich2023family, deletion, duplication, and swapping of random words) to the text $t$ with each technique applied 6 times. The modality data $m$ remains unchanged.
@@
where $E$ is intfloat/e5-mistral-7b-instruct [^21] https://hf.co/intfloat/e5-mistral-7b-instruct. It used to be SoTA on MTEB benchmark muennighoff2022mteb model.
@@ D.4 Inference and Evaluation
Appendix E LLM-as-a-judge details
E.1 Data collection
@@ E.3 Model selection
Encoders
RuModernBERT-base 150M 8,192 1800 0.964 $\pm$ 0.002 0.975 0.955 0.940 0.997
@@
Decoders
pollux-judge-7b 7.61B 131072 20 0.823 $\pm$ 0.005 0.788 0.860 0.732 0.873
@@ E.4 Model deployment
Appendix F Block prompts analysis
F.1 Prompts formulations
@@ F.2 Statistical analysis
  • Engine is the inference backend used for evaluation (one of transformers (wolf-etal-2020-transformers) or vllm [^23] https://github.com/vllm-project/vllm).
@@
Appendix G Baselines Details
G.1 Model Baselines Details
@@
Phi-3.5-vision-instruct 4B 128K microsoft/Phi-3.5-vision-instruct abdin2024phi3technicalreporthighly
Phi-4-multimodal-instruct 6B 128K microsoft/Phi-4-multimodal-instruct microsoft2025phi4minitechnicalreportcompact
Qwen2.5-Omni-3B 6B 32K Qwen/Qwen2.5-Omni-3B xu2025qwen25omnitechnicalreport
@@
Qwen2-VL-2B-Instruct 2B 32K Qwen/Qwen2-VL-2B-Instruct wang2024qwen2vlenhancingvisionlanguagemodels
@@
Qwen3-VL-2B-Instruct 2B 262K Qwen/Qwen3-VL-2B-Instruct yang2025qwen3technicalreport
@@
Qwen2.5-VL-3B-Instruct 3B 128K Qwen/Qwen2.5-VL-3B-Instruct bai2025qwen25vltechnicalreport
@@
gemma-3-12b-it 12B 128K google/gemma-3-12b-it gemmateam2025gemma3technicalreport
llava-1.5-13b-hf 13B 4K llava-hf/llava-1.5-13b-hf liu2024improvedllava
llava-next-72b-hf 72B 8K llava-hf/llava-next-72b-hf li2024llavanext-strong
@@
InternVL3-9B 9B 32K OpenGVLab/InternVL3-9B zhu2025internvl3
granite-vision-3.3-2b 3B 128K ibm-granite/granite-vision-3.3-2b granitevisionteam2025granitevisionlightweightopensource
SmolVLM-Instruct 2B 16K HuggingFaceTB/SmolVLM-Instruct marafioti2025smolvlmredefiningsmallefficient
MiniCPM-o-2_6 9B 32K openbmb/MiniCPM-o-2_6 yao2024minicpmvgpt4vlevelmllm
@@
Qwen2.5-Omni-3B 6B 32K Qwen/Qwen2.5-Omni-3B xu2025qwen25omnitechnicalreport
@@
Qwen2-Audio-7B-Instruct 7B 32K Qwen/Qwen2-Audio-7B-Instruct chu2024qwen2audiotechnicalreport
@@
audio-flamingo-3-hf 9B 32K nvidia/audio-flamingo-3-hf goel2025audioflamingo3advancing
SeaLLMs-Audio-7B 8B 8K SeaLLMs/SeaLLMs-Audio-7B zhang2024seallms3openfoundation
ultravox-v0_2 8B 8K fixie-ai/ultravox-v0_2 team2024ultravox
@@
LLaVA-NeXT-Video-7B-hf 7B 4K llava-hf/LLaVA-NeXT-Video-7B-hf liu2024llavanext
Qwen2-VL-2B-Instruct 2B 32K Qwen/Qwen2-VL-2B-Instruct wang2024qwen2vlenhancingvisionlanguagemodels
@@
Qwen2.5-VL-3B-Instruct 3B 128K Qwen/Qwen2.5-VL-3B-Instruct bai2025qwen25vltechnicalreport
@@
Qwen3-VL-2B-Instruct 2B 262K Qwen/Qwen3-VL-2B-Instruct yang2025qwen3technicalreport
@@
Qwen2.5-Omni-3B 6B 32K Qwen/Qwen2.5-Omni-3B xu2025qwen25omnitechnicalreport
@@
MiniCPM-o-2_6 9B 32K openbmb/MiniCPM-o-2_6 yao2024minicpmvgpt4vlevelmllm
InternVL3_5-4B 5B 40K OpenGVLab/InternVL3_5-4B wang2025internvl3
@@
InternVL3-9B-Instruct 9B 32K OpenGVLab/InternVL3-9B zhu2025internvl3
InternVL3-9B 9B 32K OpenGVLab/InternVL3-9B zhu2025internvl3
@@
GPT 4 0.431 0.29 / 0.29 0.23 / 0.5 0.31 / 0.4 0.27 / 0.77 0.39 / 0.21 0.01 / 0.17
Qwen2.5-VL-72B-Inst 0.353 0.53 / 0.32 0.21 / 0.48 0.4 / 0.49 0.43 / 0.73 0.35 / 0.16 0.02 / 0.23
Qwen2-VL-72B-Inst 0.332 0.12 / 0.14 0.1 / 0.47 0.29 / 0.55 0.14 / 0.6 0.14 / 0.19 0.01 / 0.06
Qwen2.5-VL-32B-Inst 0.242 0.47 / 0.22 0.16 / 0.36 0.33 / 0.4 0.3 / 0.7 0.26 / 0.1 0.0 / 0.12
Qwen2.5-VL-7B-Inst 0.238 0.19 / 0.19 0.11 / 0.25 0.23 / 0.3 0.33 / 0.52 0.2 / 0.26 0.02 / 0.1
llava-next-110b-hf 0.235 0.06 / 0.1 0.09 / 0.29 0.11 / 0.2 0.31 / 0.58 0.12 / 0.18 0.01 / 0.01
llava-next-72b-hf 0.226 0.09 / 0.11 0.07 / 0.28 0.09 / 0.18 0.21 / 0.49 0.11 / 0.15 0.0 / 0.0
Phi-3.5-vision-inst 0.226 0.12 / 0.17 0.03 / 0.09 0.08 / 0.15 0.21 / 0.4 0.19 / 0.24 0.01 / 0.02
Qwen2.5-Omni-7B 0.211 0.11 / 0.15 0.08 / 0.29 0.13 / 0.29 0.23 / 0.49 0.14 / 0.18 0.03 / 0.08
Qwen2-VL-7B-Inst 0.195 0.01 / 0.13 0.05 / 0.36 0.1 / 0.34 0.19 / 0.5 0.02 / 0.16 0.01 / 0.02
SmolVLM-Inst 0.189 0.01 / 0.16 0.03 / 0.21 0.11 / 0.33 0.14 / 0.51 0.14 / 0.22 0.02 / 0.03
Qwen3-VL-8B-Inst 0.184 0.0 / 0.07 0.11 / 0.43 0.14 / 0.32 0.23 / 0.58 0.05 / 0.15 0.04 / 0.1
Phi-4-multimodal-inst 0.183 0.26 / 0.34 0.11 / 0.28 0.09 / 0.19 0.07 / 0.35 0.04 / 0.08 0.04 / 0.07
MiniCPM-o-2_6 0.177 0.04 / 0.06 0.08 / 0.29 0.15 / 0.31 0.26 / 0.49 0.06 / 0.09 0.0 / 0.0
Qwen2.5-Omni-3B 0.172 0.04 / 0.12 0.07 / 0.21 0.16 / 0.26 0.22 / 0.45 0.04 / 0.19 0.01 / 0.04
InternVL3-9B 0.169 0.01 / 0.04 0.04 / 0.26 0.19 / 0.37 0.09 / 0.44 0.0 / 0.07 0.01 / 0.02
Qwen2-VL-2B-Inst 0.165 0.0 / 0.09 0.06 / 0.28 0.11 / 0.32 0.12 / 0.44 0.03 / 0.09 0.04 / 0.05
gemma-3-27b-it 0.146 0.0 / 0.01 0.05 / 0.38 0.04 / 0.17 0.07 / 0.44 0.0 / 0.02 0.01 / 0.04
granite-vision-3.3-2b 0.142 0.0 / 0.08 0.01 / 0.1 0.0 / 0.08 0.11 / 0.2 0.08 / 0.19 0.02 / 0.02
Qwen2.5-VL-3B-Inst 0.138 0.02 / 0.1 0.03 / 0.12 0.13 / 0.25 0.15 / 0.3 0.03 / 0.16 0.0 / 0.04
Qwen3-VL-2B-Inst 0.124 0.0 / 0.03 0.06 / 0.34 0.11 / 0.26 0.14 / 0.47 0.05 / 0.08 0.01 / 0.04
llava-1.5-13b-hf 0.113 0.01 / 0.14 0.01 / 0.1 0.01 / 0.08 0.1 / 0.32 0.02 / 0.25 0.01 / 0.0
@@
GPT 4 0.431 0.64 / 0.55 0.59 / 0.54 0.1 / 0.4 0.69 / 0.78 0.7 / 0.63
Qwen2.5-VL-72B-Inst 0.353 0.01 / 0.05 0.24 / 0.28 0.11 / 0.25 0.65 / 0.66 0.63 / 0.55
Qwen2-VL-72B-Inst 0.332 0.25 / 0.4 0.54 / 0.67 0.18 / 0.33 0.31 / 0.48 0.65 / 0.69
Qwen2.5-VL-32B-Inst 0.242 0.0 / 0.03 0.0 / 0.03 0.01 / 0.09 0.47 / 0.51 0.43 / 0.31
Qwen2.5-VL-7B-Inst 0.238 0.13 / 0.17 0.16 / 0.3 0.06 / 0.13 0.36 / 0.56 0.24 / 0.45
llava-next-110b-hf 0.235 0.25 / 0.34 0.25 / 0.45 0.06 / 0.13 0.24 / 0.33 0.54 / 0.55
llava-next-72b-hf 0.226 0.19 / 0.27 0.4 / 0.47 0.07 / 0.13 0.27 / 0.33 0.52 / 0.55
Phi-3.5-vision-inst 0.226 0.32 / 0.53 0.22 / 0.32 0.07 / 0.11 0.41 / 0.6 0.33 / 0.34
Qwen2.5-Omni-7B 0.211 0.09 / 0.21 0.12 / 0.22 0.06 / 0.12 0.3 / 0.56 0.28 / 0.46
Qwen2-VL-7B-Inst 0.195 0.04 / 0.38 0.03 / 0.52 0.09 / 0.19 0.08 / 0.38 0.11 / 0.58
SmolVLM-Inst 0.189 0.19 / 0.36 0.11 / 0.27 0.06 / 0.09 0.36 / 0.44 0.14 / 0.24
Qwen3-VL-8B-Inst 0.184 0.03 / 0.08 0.14 / 0.19 0.1 / 0.22 0.08 / 0.17 0.39 / 0.41
Phi-4-multimodal-inst 0.183 0.05 / 0.12 0.3 / 0.4 0.08 / 0.15 0.02 / 0.07 0.44 / 0.46
MiniCPM-o-2_6 0.177 0.07 / 0.19 0.2 / 0.32 0.01 / 0.1 0.2 / 0.24 0.34 / 0.4
Qwen2.5-Omni-3B 0.172 0.07 / 0.22 0.09 / 0.22 0.06 / 0.12 0.14 / 0.58 0.13 / 0.35
InternVL3-9B 0.169 0.17 / 0.24 0.37 / 0.4 0.06 / 0.16 0.01 / 0.16 0.21 / 0.4
Qwen2-VL-2B-Inst 0.165 0.04 / 0.32 0.04 / 0.38 0.08 / 0.14 0.09 / 0.31 0.16 / 0.44
gemma-3-27b-it 0.146 0.05 / 0.1 0.49 / 0.46 0.04 / 0.27 0.0 / 0.12 0.21 / 0.24
granite-vision-3.3-2b 0.142 0.27 / 0.38 0.08 / 0.22 0.04 / 0.07 0.2 / 0.52 0.2 / 0.26
Qwen2.5-VL-3B-Inst 0.138 0.02 / 0.26 0.04 / 0.21 0.05 / 0.11 0.09 / 0.47 0.08 / 0.35
Qwen3-VL-2B-Inst 0.124 0.01 / 0.04 0.14 / 0.19 0.05 / 0.11 0.03 / 0.1 0.21 / 0.25
llava-1.5-13b-hf 0.113 0.01 / 0.34 0.01 / 0.25 0.0 / 0.09 0.01 / 0.48 0.03 / 0.22
@@
Human Baseline 0,895 0.98 0.95 0.75 0.91
Qwen/Qwen3-Omni-30B-A3B-Instruct 0.531 0.69 / 0.66 0.7 / 0.72 0.43 / 0.37 0.39 / 0.28
Qwen/Qwen2.5-Omni-7B 0.464 0.55 / 0.64 0.55 / 0.65 0.36 / 0.4 0.37 / 0.18
Qwen/Qwen2.5-Omni-3B 0.377 0.41 / 0.57 0.36 / 0.64 0.3 / 0.35 0.35 / 0.04
fixie-ai/ultravox-v0_4 0.268 0.3 / 0.28 0.34 / 0.31 0.29 / 0.24 0.3 / 0.09
fixie-ai/ultravox-v0_5-llama-3_1-8b 0.266 0.3 / 0.26 0.34 / 0.24 0.29 / 0.24 0.31 / 0.15
fixie-ai/ultravox-v0_4_1-llama-3_1-8b 0.264 0.29 / 0.28 0.35 / 0.24 0.31 / 0.24 0.28 / 0.11
nvidia/audio-flamingo-3-hf 0.259 0.21 / 0.42 0.41 / 0.59 0.15 / 0.28 0.0 / 0.0
fixie-ai/ultravox-v0_4_1-mistral-nemo 0.257 0.18 / 0.32 0.27 / 0.39 0.18 / 0.33 0.26 / 0.14
fixie-ai/ultravox-v0_6-llama-3_1-8b 0.247 0.29 / 0.21 0.26 / 0.21 0.3 / 0.23 0.31 / 0.16
fixie-ai/ultravox-v0_3 0.242 0.29 / 0.26 0.35 / 0.3 0.27 / 0.22 0.22 / 0.04
Qwen/Qwen2-Audio-7B-Instruct 0.223 0.18 / 0.43 0.15 / 0.52 0.17 / 0.28 0.05 / 0.0
fixie-ai/ultravox-v0_6-qwen-3-32b 0.173 0.39 / 0.04 0.37 / 0.02 0.47 / 0.09 0.0 / 0.0
fixie-ai/ultravox-v0_2 0.139 0.01 / 0.24 0.0 / 0.38 0.01 / 0.25 0.23 / 0.0
Qwen/Qwen-Audio-Chat 0.116 0.01 / 0.27 0.01 / 0.39 0.01 / 0.23 0.0 / 0.0
SeaLLMs/SeaLLMs-Audio-7B 0.103 0.08 / 0.14 0.02 / 0.2 0.14 / 0.23 0.0 / 0.01
fixie-ai/ultravox-v0_5-llama-3_2-1b 0.102 0.04 / 0.23 0.08 / 0.26 0.05 / 0.16 0.0 / 0.0
fixie-ai/ultravox-v0_3-llama-3_2-1b 0.1 0.06 / 0.21 0.08 / 0.27 0.05 / 0.13 0.0 / 0.0

### v3

@@ B.3.4 Quantitative reasoning
Static counting I A V
Count entities in a single frame/image. Complementary to pattern recognition skills that focus on repeated structures.
Temporal counting A V
Count events and objects over time in audio/video streams. Complementary to temporal pattern recognition skills that focus on repeated structures.
@@
Solve math problems grounded in visual contexts (diagrams, plots, tables, scenes) or multimodal inputs. MathVista Lu et al. (2024) is explicitly proposed to evaluate mathematical reasoning of foundation models in visual contexts.
Mathematical reasoning I A V
B.3.5 Other reasoning
Critical thinking I A V
Evaluate evidence, reconcile inconsistencies, and make defensible judgments under ambiguous multimodal context.
Counterfactual robustness I A V
Maintain correct reasoning when interventions are applied (e.g. answer option in MCQ do not contain the correct answer, or the question relates to an entity not present in the image).
Problem decomposition I A V
Solve by breaking into intermediate steps; multi-step chains across modalities. ScienceQA Lu et al. (2022) is explicitly proposed to evaluate problem decomposition skills of foundation models.
Comparative reasoning I A V
Answer questions requiring comparisons (more/less, same/different) across instances or across time.
Appendix C Data leakage details
@@
C.2 Neighbor Generation and Feature Extraction.
For each original data point $(t,m)$ we generate $K=24$ perturbed "neighbors". We apply four distinct perturbation techniques (masking and predicting the masks with Fred-T5 model [^20] https://hf.co/ai-forever/FRED-T5-1.7B, Zmitrovich et al. (2024), deletion, duplication, and swapping of random words) to the text $t$ with each technique applied 6 times. The modality data $m$ remains unchanged.
@@
where $E$ is intfloat/e5-mistral-7b-instruct [^21] https://hf.co/intfloat/e5-mistral-7b-instruct. It used to be SoTA on MTEB benchmark Muennighoff et al. (2022) model.
@@ C.4 Inference and Evaluation
Appendix D LLM-as-a-judge details
D.1 Data collection
@@ D.3 Model selection
Encoders RuModernBERT-base 150M 8,192 1800 0.964 $\pm$ 0.002 0.975 0.955 0.940 0.997
@@
Decoders pollux-judge-7b 7.61B 131072 20 0.823 $\pm$ 0.005 0.788 0.860 0.732 0.873
@@ D.4 Model deployment
D.5 Model analysis
To validate our LLM-as-a-judge, we performed a sanity check demonstrating high heuristic consistency: the model aligns with human labels in 99.6% of cases where Exact Match (EM) is 1, and 97.7% when EM is 0 but humans deem the response correct. This reliability extends across task types, with agreement rates of 98.7% for multiple-choice and 96% for free-form generation. Statistical analysis confirms the judge is resilient to common biases, showing near-zero Pearson and Spearman correlations with answer length ($r\approx 0.001$, $\rho\approx-0.008$) and gold label position ($r/\rho\approx-0.0005$). Qualitative analysis reveals that errors primarily occur in tasks involving complex LaTeX formatting or multi-step reasoning. Since our model is designed as a binary classifier without a rationale-generation component, it occasionally struggles with the ambiguous justifications or symbolic intricacies inherent in these specialized domains.
Appendix E Block prompts analysis
E.1 Prompts formulations
@@ E.2 Statistical analysis
  • Engine is the inference backend used for evaluation (one of transformers (Wolf et al., 2020) or vllm [^23] https://github.com/vllm-project/vllm).
@@
Appendix F Baselines Details
F.1 Model Baselines Details
@@
Phi-3.5-vision-instruct 4B 128K microsoft/Phi-3.5-vision-instruct Abdin et al. (2024)
Phi-4-multimodal-instruct 6B 128K microsoft/Phi-4-multimodal-instruct Microsoft et al. (2025)
Qwen2.5-Omni-3B 6B 32K Qwen/Qwen2.5-Omni-3B Xu et al. (2025)
@@
Qwen2-VL-2B-Instruct 2B 32K Qwen/Qwen2-VL-2B-Instruct Wang et al. (2024a)
@@
Qwen3-VL-2B-Instruct 2B 262K Qwen/Qwen3-VL-2B-Instruct Yang et al. (2025)
@@
Qwen2.5-VL-3B-Instruct 3B 128K Qwen/Qwen2.5-VL-3B-Instruct Bai et al. (2025)
@@
gemma-3-12b-it 12B 128K google/gemma-3-12b-it Team et al. (2025a)
llava-1.5-13b-hf 13B 4K llava-hf/llava-1.5-13b-hf Liu et al. (2024b)
llava-next-72b-hf 72B 8K llava-hf/llava-next-72b-hf Li et al. (2024)
@@
InternVL3-9B 9B 32K OpenGVLab/InternVL3-9B Zhu et al. (2025)
granite-vision-3.3-2b 3B 128K ibm-granite/granite-vision-3.3-2b Team et al. (2025b)
SmolVLM-Instruct 2B 16K HuggingFaceTB/SmolVLM-Instruct Marafioti et al. (2025)
MiniCPM-o-2_6 9B 32K openbmb/MiniCPM-o-2_6 Yao et al. (2024)
@@
Qwen2.5-Omni-3B 6B 32K Qwen/Qwen2.5-Omni-3B Xu et al. (2025)
@@
Qwen2-Audio-7B-Instruct 7B 32K Qwen/Qwen2-Audio-7B-Instruct Chu et al. (2024)
@@
audio-flamingo-3-hf 9B 32K nvidia/audio-flamingo-3-hf Goel et al. (2025)
SeaLLMs-Audio-7B 8B 8K SeaLLMs/SeaLLMs-Audio-7B Zhang et al. (2024)
ultravox-v0_2 8B 8K fixie-ai/ultravox-v0_2 Team (2024)
@@
LLaVA-NeXT-Video-7B-hf 7B 4K llava-hf/LLaVA-NeXT-Video-7B-hf Liu et al. (2024c)
Qwen2-VL-2B-Instruct 2B 32K Qwen/Qwen2-VL-2B-Instruct Wang et al. (2024a)
@@
Qwen2.5-VL-3B-Instruct 3B 128K Qwen/Qwen2.5-VL-3B-Instruct Bai et al. (2025)
@@
Qwen3-VL-2B-Instruct 2B 262K Qwen/Qwen3-VL-2B-Instruct Yang et al. (2025)
@@
Qwen2.5-Omni-3B 6B 32K Qwen/Qwen2.5-Omni-3B Xu et al. (2025)
@@
MiniCPM-o-2_6 9B 32K openbmb/MiniCPM-o-2_6 Yao et al. (2024)
InternVL3_5-4B 5B 40K OpenGVLab/InternVL3_5-4B Wang et al. (2025b)
@@
InternVL3-9B-Instruct 9B 32K OpenGVLab/InternVL3-9B Zhu et al. (2025)
InternVL3-9B 9B 32K OpenGVLab/InternVL3-9B Zhu et al. (2025)
@@
Qwen3-Omni-30B-A3B-Inst 0.55 0.56 / 0.62 0.31 / 0.67 0.50 / 0.63 0.56 / 0.86 0.45 / 0.57 0.01 / 0.30
GPT 4 0.48 0.29 / 0.29 0.23 / 0.70 0.31 / 0.53 0.27 / 0.85 0.39 / 0.49 0.01 / 0.17
Qwen2.5-VL-72B-Inst 0.41 0.53 / 0.60 0.21 / 0.65 0.40 / 0.58 0.43 / 0.80 0.35 / 0.44 0.02 / 0.23
Qwen2-VL-72B-Inst 0.33 0.12 / 0.14 0.10 / 0.49 0.29 / 0.55 0.14 / 0.60 0.14 / 0.19 0.01 / 0.06
Qwen2.5-VL-32B-Inst 0.31 0.47 / 0.57 0.16 / 0.50 0.33 / 0.58 0.30 / 0.76 0.26 / 0.40 0.00 / 0.12
Qwen2.5-VL-7B-Inst 0.26 0.19 / 0.27 0.11 / 0.32 0.23 / 0.34 0.33 / 0.56 0.20 / 0.29 0.02 / 0.10
llava-next-110b-hf 0.24 0.06 / 0.10 0.09 / 0.29 0.11 / 0.20 0.31 / 0.58 0.12 / 0.18 0.01 / 0.01
Phi-3.5-vision-inst 0.23 0.12 / 0.18 0.03 / 0.10 0.08 / 0.16 0.21 / 0.40 0.19 / 0.26 0.01 / 0.02
llava-next-72b-hf 0.23 0.09 / 0.11 0.07 / 0.29 0.09 / 0.19 0.21 / 0.49 0.11 / 0.15 0.00 / 0.00
Qwen2.5-Omni-7B 0.23 0.11 / 0.18 0.08 / 0.35 0.13 / 0.32 0.23 / 0.53 0.14 / 0.23 0.03 / 0.08
Qwen2-VL-7B-Inst 0.20 0.01 / 0.13 0.05 / 0.36 0.10 / 0.34 0.19 / 0.50 0.02 / 0.16 0.01 / 0.02
SmolVLM-Inst 0.19 0.01 / 0.16 0.03 / 0.23 0.11 / 0.34 0.14 / 0.52 0.14 / 0.22 0.02 / 0.04
Qwen3-VL-8B-Inst 0.19 0.00 / 0.07 0.11 / 0.45 0.14 / 0.33 0.23 / 0.59 0.05 / 0.15 0.04 / 0.10
Phi-4-multimodal-inst 0.18 0.26 / 0.35 0.11 / 0.30 0.09 / 0.20 0.07 / 0.36 0.04 / 0.08 0.04 / 0.07
MiniCPM-o-2_6 0.18 0.04 / 0.06 0.08 / 0.31 0.15 / 0.33 0.26 / 0.50 0.06 / 0.09 0.00 / 0.00
Qwen2.5-Omni-3B 0.18 0.04 / 0.14 0.07 / 0.25 0.16 / 0.29 0.22 / 0.48 0.04 / 0.21 0.01 / 0.04
InternVL3-9B 0.17 0.01 / 0.04 0.04 / 0.27 0.19 / 0.37 0.09 / 0.45 0.00 / 0.07 0.01 / 0.02
Qwen2-VL-2B-Inst 0.17 0.00 / 0.09 0.06 / 0.28 0.11 / 0.32 0.12 / 0.44 0.03 / 0.09 0.04 / 0.05
gemma-3-27b-it 0.15 0.00 / 0.01 0.05 / 0.38 0.04 / 0.17 0.07 / 0.44 0.00 / 0.02 0.01 / 0.04
granite-vision-3.3-2b 0.14 0.00 / 0.08 0.01 / 0.10 0.00 / 0.08 0.11 / 0.20 0.08 / 0.20 0.02 / 0.02
Qwen2.5-VL-3B-Inst 0.14 0.02 / 0.11 0.03 / 0.13 0.13 / 0.26 0.15 / 0.32 0.03 / 0.18 0.00 / 0.04
Qwen3-VL-2B-Inst 0.12 0.00 / 0.04 0.06 / 0.35 0.11 / 0.27 0.14 / 0.47 0.05 / 0.08 0.01 / 0.04
llava-1.5-13b-hf 0.12 0.01 / 0.16 0.01 / 0.10 0.01 / 0.08 0.10 / 0.33 0.02 / 0.26 0.01 / 0.01
@@
Qwen3-Omni-30B-A3B-Inst 0.55 0.77 / 0.85 0.64 / 0.70 0.11 / 0.34 0.70 / 0.78 0.63 / 0.64
GPT 4 0.48 0.64 / 0.69 0.59 / 0.66 0.10 / 0.40 0.69 / 0.79 0.70 / 0.73
Qwen2.5-VL-72B-Inst 0.41 0.01 / 0.05 0.24 / 0.32 0.11 / 0.25 0.65 / 0.76 0.63 / 0.69
Qwen2-VL-72B-Inst 0.33 0.25 / 0.40 0.54 / 0.68 0.18 / 0.33 0.31 / 0.48 0.65 / 0.69
Qwen2.5-VL-32B-Inst 0.31 0.00 / 0.03 0.00 / 0.04 0.01 / 0.09 0.47 / 0.77 0.43 / 0.62
Qwen2.5-VL-7B-Inst 0.26 0.13 / 0.17 0.16 / 0.32 0.06 / 0.13 0.36 / 0.62 0.24 / 0.47
llava-next-110b-hf 0.24 0.25 / 0.34 0.25 / 0.45 0.06 / 0.13 0.24 / 0.33 0.54 / 0.55
Phi-3.5-vision-inst 0.23 0.32 / 0.53 0.22 / 0.33 0.07 / 0.11 0.41 / 0.60 0.33 / 0.35
llava-next-72b-hf 0.23 0.19 / 0.27 0.40 / 0.47 0.07 / 0.13 0.27 / 0.33 0.52 / 0.55
Qwen2.5-Omni-7B 0.23 0.09 / 0.23 0.12 / 0.24 0.06 / 0.12 0.30 / 0.60 0.28 / 0.48
Qwen2-VL-7B-Inst 0.20 0.04 / 0.38 0.03 / 0.52 0.09 / 0.19 0.08 / 0.38 0.11 / 0.58
SmolVLM-Inst 0.19 0.19 / 0.36 0.11 / 0.27 0.06 / 0.10 0.36 / 0.44 0.14 / 0.24
Qwen3-VL-8B-Inst 0.19 0.03 / 0.08 0.14 / 0.20 0.10 / 0.22 0.08 / 0.17 0.39 / 0.41
Phi-4-multimodal-inst 0.18 0.05 / 0.12 0.30 / 0.40 0.08 / 0.15 0.02 / 0.07 0.44 / 0.46
MiniCPM-o-2_6 0.18 0.07 / 0.19 0.20 / 0.35 0.01 / 0.10 0.20 / 0.24 0.34 / 0.41
Qwen2.5-Omni-3B 0.18 0.07 / 0.23 0.09 / 0.24 0.06 / 0.12 0.14 / 0.60 0.13 / 0.36
InternVL3-9B 0.17 0.17 / 0.25 0.37 / 0.44 0.06 / 0.16 0.01 / 0.16 0.21 / 0.40
Qwen2-VL-2B-Inst 0.17 0.04 / 0.32 0.04 / 0.38 0.08 / 0.14 0.09 / 0.31 0.16 / 0.44
gemma-3-27b-it 0.15 0.05 / 0.12 0.49 / 0.55 0.04 / 0.28 0.00 / 0.12 0.21 / 0.25
granite-vision-3.3-2b 0.14 0.27 / 0.38 0.08 / 0.22 0.04 / 0.07 0.20 / 0.52 0.20 / 0.26
Qwen2.5-VL-3B-Inst 0.14 0.02 / 0.26 0.04 / 0.22 0.05 / 0.11 0.09 / 0.50 0.08 / 0.35
Qwen3-VL-2B-Inst 0.12 0.01 / 0.04 0.14 / 0.19 0.05 / 0.11 0.03 / 0.10 0.21 / 0.25
llava-1.5-13b-hf 0.12 0.01 / 0.37 0.01 / 0.25 0.00 / 0.09 0.01 / 0.49 0.03 / 0.22
@@
Human Baseline 0.895 0.98 0.95 0.75 0.91
Qwen3-Omni-30B-A3B-Instruct 0.56 0.69 / 0.77 0.70 / 0.78 0.43 / 0.44 0.39 / 0.28
Qwen2.5-Omni-7B 0.47 0.55 / 0.67 0.55 / 0.70 0.36 / 0.41 0.37 / 0.18
Qwen2.5-Omni-3B 0.38 0.41 / 0.58 0.36 / 0.64 0.30 / 0.36 0.35 / 0.04
MiniCPM-o-2_6 0.37 0.40 / 0.55 0.47 / 0.64 0.25 / 0.32 0.31 / 0.01
ultravox-v0_6-qwen-3-32b 0.32 0.39 / 0.42 0.37 / 0.40 0.47 / 0.51 0.00 / 0.00
ultravox-v0_5-llama-3_1-8b 0.31 0.30 / 0.37 0.34 / 0.41 0.29 / 0.31 0.31 / 0.15
ultravox-v0_4_1-llama-3_1-8b 0.31 0.29 / 0.36 0.35 / 0.42 0.31 / 0.32 0.28 / 0.11
ultravox-v0_4 0.30 0.30 / 0.38 0.34 / 0.43 0.29 / 0.31 0.30 / 0.09
ultravox-v0_6-llama-3_1-8b 0.30 0.29 / 0.36 0.26 / 0.40 0.30 / 0.32 0.31 / 0.16
ultravox-v0_3 0.28 0.29 / 0.35 0.35 / 0.43 0.27 / 0.29 0.22 / 0.04
ultravox-v0_4_1-mistral-nemo 0.26 0.18 / 0.34 0.27 / 0.42 0.18 / 0.34 0.26 / 0.14
audio-flamingo-3-hf 0.26 0.21 / 0.42 0.41 / 0.59 0.15 / 0.28 0.00 / 0.00
Qwen2-Audio-7B-Instruct 0.22 0.18 / 0.43 0.15 / 0.52 0.17 / 0.28 0.05 / 0.00
ultravox-v0_2 0.14 0.01 / 0.25 0.00 / 0.38 0.01 / 0.25 0.23 / 0.00
ultravox-v0_5-llama-3_2-1b 0.12 0.04 / 0.25 0.08 / 0.29 0.05 / 0.23 0.00 / 0.00
Qwen-Audio-Chat 0.12 0.01 / 0.28 0.01 / 0.40 0.01 / 0.24 0.00 / 0.00
ultravox-v0_3-llama-3_2-1b 0.12 0.06 / 0.24 0.08 / 0.29 0.05 / 0.20 0.00 / 0.00
SeaLLMs-Audio-7B 0.10 0.08 / 0.14 0.02 / 0.20 0.14 / 0.23 0.00 / 0.01